OpenStudy (anonymous):

how would i add this up? (3x^3y)(3x^3y)(-4x^2y^4) (-4x^2y^4) (-4x^2y^4)

OpenStudy (pfenn1):

Is this your expression?\[(3x^3y)(3x^3y)(-4x^2y^4) (-4x^2y^4) (-4x^2y^4)\]

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

well, I would put like terms together so\[(3x^3y)(3x^3y)(-4x^2y^4) (-4x^2y^4) (-4x^2y^4)=576 x ^{3+3+2+2+2}y ^{1+1+4+4+4}=?\]

OpenStudy (pfenn1):

sorry\[=-576 x ^{3+3+2+2+2}y ^{1+1+4+4+4}=?\]

OpenStudy (anonymous):

^12y ^12

OpenStudy (pfenn1):

I think you didn't get that quite right. A typo maybe?